A lottery winner claims a prize of $1.2 million, payable over 30 years at $40,000 per year. If the first payment is made immediately, what is this prize really worth given the annual rate of 6.8%?
Answer #1

We can use the present value of annuity due formula to find the answer:

\small \textup{PVA} = \textup{A}\left [\frac{(1-(1+i)^{-n+1})}{i} \right ] + \textup{A}

Where,
PVA = Present value of the annuity
A = Annuity
i = Interest rate in decimal form (i.e 6.8% = 0.068)
n = Number of years

Therefore,

\small \textup{PVA} = \textup{40000}\left [\frac{(1-(1+0.068)^{-30+1})}{0.068} \right ] + \textup{40000}

\small = \textup{40000}\left [\frac{(1-(1.068)^{-29})}{0.068} \right ] + \textup{40000}

\small = \textup{40000}\left [\frac{(1-0.14839992)}{0.068} \right ] + \textup{40000}

\small = \textup{40000}\left [\frac{0.85160008}{0.068} \right ] + \textup{40000}

\small = (\textup{40000}\times 12.5235306)+ \textup{40000}

\small = \textup{500,941.22}+ \textup{40000}

\small = \$$540,941.22

Therefore, the prize is worth $540,941.22 today.